# The Dynamical History of the Kepler-221 Planet System

### v1

@@ 4.4 Success Rate Analysis
To understand the success of resonance reformation post-collision, we run 200 simulations with different parameters (see Table 3). In 101 of them planets $\mathrm{d_{1}}$ and $\mathrm{d_{2}}$ collide, as shown in Figure 4, while in the other simulations planets c and $\mathrm{d_{1}}$ collide, or some planets are ejected out of the system. The overall success rate in post-collision phase is around 10% (9 successful in 101 simulations). In the simulation, an outward torque is applied on planet b corresponding to a mass loss of 2% to promote the convergent migration between planets. Tidal damping is applied on all planets with the tidal $Q_{\mathrm{phy}}$ parameter the same on each planet and varies from 3 to 10. A simulation is categorized as a successful simulation if planet b/c/e successfully gets trapped in the 0th-order 3BR.
@@
The resonance reformation process is a critical step of the model. To obtain meaningful statistics on its viability we conduct a parameter study of Phases II and III. Altogether 200 simulations are run varying parameters such as tidal damping strength and additional torque on planet b. As it is already shown that convergent migration is critical, an outward torque is applied on planet b with $t^{\mathrm{b}}_{\mathrm{0,III}}$ between $[2.5,7.5]\,\mathrm{Myr}$ and $t^{\mathrm{b}}_{\mathrm{a,III}}$ in the range of $[6\times 10^{7},2\times 10^{8}]\mathrm{yr}$ (see above, Section 4.3). Tidal damping is applied on all planets with the tidal $Q_{\mathrm{phy}}$ parameter the same for each planet and varying from 3 to 10. A simulation is categorized as successful if planet b/c/e successfully gets trapped in the 0th-order 3BR. Based on these simulations, we identify milestones that must be taken to guarantee resonance reformation. The results are presented graphically in Figure 4 in the form of a decision tree. Although the bulk success rate is low, less than 10%, Figure 4 shows that it can be significant when certain conditions and Ansatzes are met.
@@
Reformation of the 3BR is only possible when $P_{\mathrm{c}}/P_{\mathrm{b}}>2$ and $P_{\mathrm{e}}/P_{\mathrm{c}}>3$ immediately post-collision (this happens in 40% of the simulations). Here, the period ratio is measured as the average in the first $100\,\mathrm{yr}$ after merging of planets $\mathrm{d_{1}}$ and $\mathrm{d_{2}}$. Period ratios in excess of exact resonance are needed in order to set up conditions for convergently migrating planets into resonance (see Section 4.2). In 68% of these simulations planets b and c undergo a relatively slow convergent migration ($-0.018\mathrm{Myr}^{-1}<\Delta(P_{\mathrm{c}}/P_{\mathrm{b}})/\Delta t<0$) and the b/c 2BR successfully reforms. Due to the outward torque on b and the tidal damping on the planets, planet c would migrate outward and undergo convergent migration with planet e after the formation of b/c 2BR. The trapping of planets b/c/e prefer slow migration of c (with $-0.015\mathrm{Myr}^{-1}<\Delta(P_{\mathrm{e}}/P_{\mathrm{c}})/\Delta t<0$, occurring in 78% of the simulation). Otherwise, if $P_{\mathrm{e}}/P_{\mathrm{c}}$ decreases too rapidly, the period ratio between planets c and e will quickly drop below 3 and the trapping of c/e 2BR and b/c/e 3BR becomes impossible. When all these migration criteria are satisfied, the b/c/e 3BR has a chance of forming in 43% of the simulation (9 out of 21).
@@
In reality, the situation could belong into this category because we omitted the higher $Q_{\mathrm{phy}}$ parameters from our parameter study due to computational constraints. Specifically, we have enforced a maximum simulation time of $10\,\mathrm{Myr}$, within which the resonance reformation must take place. Therefore, a more gentle convergent migration during the formation of b/c 2BR and b/c/e 3BR would naturally satisfy conditions 3) and 4) in Figure 4 with a higher tidal Q parameter (Wu, 2005; Jackson et al., 2008). On the other hand, it is impossible to reform the b/c/e 3BR with $Q_{\mathrm{phy}}<3$ due to too rapid migration. Therefore, given the Ansatz that a collision took place, the (maximum) success rate of the b/c/e resonance reformation stands at about 17% (=$0.4{\times}0.43$).
@@
Finally, we run a similar success rate analysis for mass model M2d in Table 2 assuming planet d is about twice as massive as other planets. The success rate of 3BR reformation in Phase III is around 10%, similar to mass model M2a. This implies that a massive planet d, potentially due to the merging of planets d1 and d2, has minimal impact on the reformation of b/c/e 3BR in Phase III. In conclusion, the reformation of 3BR in the post-collision phase appears viable under plausible physical conditions.
@@ 5 Orbital Expansion (Phase IV)
After the formation of b/c/e 3BR, discussed in Section 4, the period ratios between b, c, and e are still close to integer. On the other hand, planets b, c, and e in the Kepler-221 system are currently locked into three-body resonance, relatively far away from integer period ratio, with $P_{\mathrm{c}}/P_{\mathrm{b}}=2.035$ and $P_{\mathrm{e}}/P_{\mathrm{c}}=3.228$. This implies that Phase IV – the orbital expansion phase – must have happened in the history of the Kepler-221 system (Goldberg & Batygin, 2021), regardless of the formation origin of the 3BR. The 3BR expansion excluding planet d in the Kepler-221 system was already investigated in Goldberg & Batygin (2021). In this section, we run the expansion phase including planet d, and demonstrate under which conditions tidal damping acting on the b, c, and e three-body resonance is capable of expanding these planets to their current observed period ratios. We will show that planet d can significantly affect the orbital expansion phase due to b/c/d 3BR crossing, which could put constraints on the system parameters including the mass ratio of planets.
@@ 5.1 Orbital Expansion: Failure and Success
Figure 6: Possible mechanisms which ensure successful orbital expansion of the Kepler-221 system. Two simulations with identical initial positions of the planets and different tidal strengths, torque on planet b, and planet masses are shown from top to bottom. (i) Top panels: A temporal outward torque on planet b with masses according to M1 in Table 2; (ii) Bottom panels: different initial masses of planets according to M2a in Table 2.
@@
In the orbital expansion phase, a successful simulation matching the observed period ratio has two requirements: i) planets b/c/e should stay in resonance and expand to the current period ratio $P_{\mathrm{c}}/P_{\mathrm{b}}=2.035$ and $P_{\mathrm{e}}/P_{\mathrm{c}}=3.228$; ii) when planets b/c/e reach the current period ratio, planet d should also match the observed period ratio with $P_{\mathrm{d}}/P_{\mathrm{c}}=1.765$ and $P_{\mathrm{e}}/P_{\mathrm{d}}=1.829$. A successful expansion matching these two conditions like in panels d and g in Figure 2 is not guaranteed.
@@
Because the orbital expansion phase is not related to the formation process of b/c/e 3BR, we use an optimized model to form the b/c/e 3BR to simplify the parameter setup like the initial masses of the planets. We initialize five planets (b, c, $\mathrm{d_{1}}$, $\mathrm{d_{2}}$ and e) in the 1st-order resonance chain consistent with Phase I and adiabatically decrease the masses of $\mathrm{d_{1}}$ and $\mathrm{d_{2}}$ to zero to ensure that the remaining planet b, c and e are still in 3BR close to integer period ratio. Then we add planet d back. The detailed setup of the optimized model is discussed in Appendix B.
@@
Figure 5 shows the outcome of two unsuccessful simulations with the same masses of planets following M1 in Table 2 and different assumptions on the initial value of $P_{\mathrm{d}}/P_{\mathrm{c}}$. Initially, the planets are close to their corresponding 2:1 and 3:1 period ratios, which is a condition for the trapping in the 0th-order 3BR (see Section 4.2). The dashed lines in Figure 5 indicate different 3BRs characterized by $p$, $q$ and $r$ in Equation (7). Tidal dissipation moves planets along $(2,3,5)$ 0th-order 3BR between b, c, and e (Charalambous et al., 2018; Papaloizou et al., 2018), indicated in the left panels of Figure 5 by the blue dashed line.
@@
While the b/c/e system starts from near-exact resonance locations, there is some freedom to choose the initial starting point of planet d. In Figure 5, panels a and b present a simulation where planet d is positioned below the $(p,q,r)=(3,5,8)$ 3BR resonance line. In addition, only tidal damping operates on the planets and all planets have their $Q_{\mathrm{phys}}=4.6$, which is consistent with the successful simulation in Figure 2. These low Q-values are solely adopted for computational expediency since they speed up the simulation. However, as will be discussed in more detail in Section 6.3 the young age demands a very efficient damping mechanism. The adopted value for $Q_{\mathrm{phys}}$ in the orbital expansion phase simulations does not affect the simulation outcome. A larger $Q_{\mathrm{phys}}$ will not change the simulation outcome, but it will make the duration of the orbital expansion to the observed period ratio longer. Therefore, the age of the system (younger than $650\;\mathrm{Myr}$) might be inconsistent with the simulation, as explained in Section 6.3. During tidal expansion the periods of planets b, c, and e change, following (by definition) the $(2,3,5)$ 3BR line (panel a), while $P_{\mathrm{d}}$ remains largely constant. As a result, the increase of $P_{\mathrm{e}}$ and decrease of $P_{\mathrm{c}}$ lead to the increase of both $P_{\mathrm{e}}/P_{\mathrm{d}}$ and $P_{\mathrm{d}}/P_{\mathrm{c}}$ and the system moves away from the (3,5,8) resonance line (panel b). Therefore, during the orbital expansion planets c, d, and e would not encounter this 3BR, which could dislodge it from the b/c/e resonance (see below), and planets b, c, and e successfully expand to the observed period ratio (panel a). Nevertheless, this scenario fails as the present location of planet d with respect to c an e (the black cross in Figure 5b) cannot be reached.
@@
The slope in the $P_{\mathrm{e}}/P_{\mathrm{d}}$ vs $P_{\mathrm{d}}/P_{\mathrm{c}}$ plane (Figure 5b) is the result of the conservation of the 3BR resonance angle and angular momentum. Specifically, without considering the potentially disturbing effects of planet d, the slope in this plane can be analytically solved (see Appendix C):
$a_{\mathrm{cde}}=\frac{\Delta({P_{\mathrm{e}}/P_{\mathrm{d}}})}{\Delta(P_{\mathrm{d}}/P_{\mathrm{c}})}=\frac{4.90m_{\mathrm{b}}+4.94m_{\mathrm{c}}}{7.27m_{\mathrm{e}}-m_{\mathrm{b}}}$ (10)
@@
A possible solution to match the low $a_{\mathrm{cde}}$ is to decrease the initial semi-major axis of planet d, and thereby $P_{\mathrm{d}}/P_{\mathrm{c}}$, in such a way that it could reach the correct observed position. One example is shown in panels c and d of Figure 5. However, this simulation is also unsuccessful, because planets c, d, and e would encounter the $(3,5,8)$ 3BR during the orbital expansion, as shown in panel d. The resonance overlapping experienced at this point would break the b, c, and e 3BR and stop the orbital expansion of these planets (Petit et al., 2020). We find that the b/c/e 3BR are dislodged in more than 95% simulations, similar to what happened in panel c.
@@
    A (temporary) slowdown in the expansion of $P_{\mathrm{d}}/P_{\mathrm{c}}$ due to outward migration of planet c, for example through a positive torque acting on planet b. Because planets b, c, and e are linked by 3BR, the outward migration of b would also lead to the increase in the semi-major axis of c and e. An example of a scenario that includes positive torques are shown in panels a and b of Figure 6. Specifically, a positive torque, with $t_{\mathrm{a0,IV}}^{\mathrm{b}}=2.1\times 10^{10}\,\mathrm{yr}$ and $t_{\Gamma}^{\mathrm{b}}=1.4\,\mathrm{Gyr}$, is added to planet b (see below, Section 5.2.1). The torque slows down the expansion of $P_{\mathrm{d}}/P_{\mathrm{c}}$ with respect to $P_{\mathrm{e}}/P_{\mathrm{d}}$. Its decaying nature results in an S-shaped trajectory in the c/d/e period ratio plane (Figure 5b). As a result, the planets successfully reach the observed period ratio near the end of the simulation (see Section 5.2.1).
@@
Goldberg & Batygin (2021) expanded b, c, and e 3BR, but planet d was neglected because its current position is far from a 3BR with any other planet. We stress, however, that planet d is crucial in the orbital expansion phase. Although planet d is far from 3BR currently, when planets c, d, and e hit a 3BR during the orbital evolution, it could break the 3BR between b, c, and e and result in a different architecture from the observation. Therefore, it is necessary to include planet d in the orbital expansion simulation. The initial, relative positions of planet d and the c/d/e 3BR are crucial for the subsequent orbital expansion of the system.

### v2

@@ 4.4 Success rate analysis
To understand the success of resonance reformation post-collision, we run 200 simulations with different parameters (see Table 3). In 101 of them planets $\mathrm{d_{1}}$ and $\mathrm{d_{2}}$ collide, as shown in Figure 4, while in the other simulations planets c and $\mathrm{d_{1}}$ collide, or some planets are ejected out of the system. The overall success rate in the post-collision phase is around 10% (9 successful in 101 simulations). In the simulation, an outward torque is applied on planet b corresponding to a mass loss of 2% to promote the convergent migration between planets. Tidal damping is applied on all planets with the tidal $Q_{\mathrm{phy}}$ parameter the same on each planet and varies from 3 to 10. A simulation is categorized as a successful simulation if planet (b, c, e) successfully gets trapped in the zeroth-order 3BR.
@@
The resonance reformation process is a critical step of the model. To obtain meaningful statistics on its viability we conduct a parameter study of Phases II and III. Altogether 200 simulations are run varying parameters such as tidal damping strength and additional torque on planet b. As it is already shown that convergent migration is critical, an outward torque is applied on planet b with $t^{\mathrm{b}}_{\mathrm{0,III}}$ between $[2.5,7.5]\,\mathrm{Myr}$ and $t^{\mathrm{b}}_{\mathrm{a,III}}$ in the range of $[6\times 10^{7},2\times 10^{8}]\mathrm{yr}$ (see above, Section 4.3). Tidal damping is applied on all planets with the tidal $Q_{\mathrm{phy}}$ parameter the same for each planet and varying from 3 to 10. A simulation is categorized as successful if planet (b, c, e) successfully gets trapped in the zeroth-order 3BR. Based on these simulations, we identify milestones that must be taken to guarantee resonance reformation. The results are presented graphically in Figure 4 in the form of a decision tree. Although the bulk success rate is low, less than 10%, Figure 4 shows that it can be significant when certain conditions and Ansatzes are met.
@@
Reformation of the 3BR is only possible when $P_{\mathrm{c}}/P_{\mathrm{b}}>2$ and $P_{\mathrm{e}}/P_{\mathrm{c}}>3$ immediately post-collision (this happens in 40% of the simulations). Here, the period ratio is measured as the average in the first $100\,\mathrm{yr}$ after merging of planets $\mathrm{d_{1}}$ and $\mathrm{d_{2}}$. Period ratios above exact resonance are needed in order to set up conditions for convergently migrating planets into resonance (see Section 4.2). In 68% of these simulations, planets b and c undergo a relatively slow convergent migration ($-0.018\mathrm{Myr}^{-1}<\Delta(P_{\mathrm{c}}/P_{\mathrm{b}})/\Delta t<0$) and the b/c 2BR successfully reforms. Due to the outward torque on b and the tidal damping on the planets, planet c would migrate outward and undergo convergent migration with planet e after the formation of b/c 2BR. The trapping of planets (b, c, e) prefer slow migration of c (with $-0.015\mathrm{Myr}^{-1}<\Delta(P_{\mathrm{e}}/P_{\mathrm{c}})/\Delta t<0$, occurring in 78% of the simulation). Otherwise, if $P_{\mathrm{e}}/P_{\mathrm{c}}$ decreases too rapidly, the period ratio between planets c and e will quickly drop below three, and the trapping of c/e 2BR and (b, c, e) 3BR becomes impossible. When all these migration criteria are satisfied, the (b, c, e) 3BR has a chance of forming in 43% of the simulation (9 out of 21).
@@
In reality, the situation could belong in this category because we omitted the higher $Q_{\mathrm{phy}}$ parameters from our parameter study due to computational constraints. Specifically, we have enforced a maximum simulation time of $10\,\mathrm{Myr}$, within which the resonance reformation must take place. Therefore, a more gentle convergent migration during the formation of b/c 2BR and (b, c, e) 3BR would naturally satisfy conditions 3) and 4) in Figure 4 with a higher tidal Q parameter (Wu, 2005; Jackson et al., 2008). On the other hand, it is impossible to reform the (b, c, e) 3BR with $Q_{\mathrm{phy}}<3$ due to too rapid migration. Therefore, given the Ansatz that a collision took place, the (maximum) success rate of the (b, c, e) resonance reformation stands at about 17% (=$0.4{\times}0.43$).
@@
Finally, we run a similar success rate analysis for mass model M2d in Table 2 assuming planet d is about twice as massive as other planets. The success rate of 3BR reformation in Phase III is around 10%, similar to mass model M2a. This implies that a massive planet d, potentially due to the merging of planets d1 and d2, has minimal impact on the reformation of (b, c, e) 3BR in Phase III. In conclusion, the reformation of 3BR in the post-collision phase appears viable under plausible physical conditions.
@@ 5 Orbital expansion (Phase IV)
After the formation of (b, c, e) 3BR, discussed in Section 4, the period ratios between b, c, and e are still close to integer. On the other hand, planets b, c, and e in the Kepler-221 system are currently locked into three-body resonance, relatively far away from integer period ratio, with $P_{\mathrm{c}}/P_{\mathrm{b}}=2.035$ and $P_{\mathrm{e}}/P_{\mathrm{c}}=3.228$. This implies that Phase IV – the orbital expansion phase – must have happened in the history of the Kepler-221 system (Goldberg & Batygin, 2021), regardless of the formation origin of the 3BR. The 3BR expansion excluding planet d in the Kepler-221 system was already investigated in Goldberg & Batygin (2021). In this section, we run the expansion phase including planet d, and demonstrate under which conditions tidal damping acting on the b, c, and e three-body resonance is capable of expanding these planets to their current observed period ratios. We will show that planet d can significantly affect the orbital expansion phase due to (b, c, d) 3BR crossing, which could put constraints on the system parameters including the mass ratio of planets.
@@ 5.1 Orbital expansion: Failure and success
Figure 6: Mechanisms that ensure successful orbital expansion of the Kepler-221 system. Two simulations with identical initial positions of the planets and different tidal strengths, torque on planet b, and planet masses are shown from top to bottom. (i) Top panels: A temporal outward torque on planet b with masses according to M1 in Table 2; (ii) Bottom panels: different initial masses of planets according to M2a in Table 2.
@@
In the orbital expansion phase, a successful simulation matching the observed period ratio has two requirements: i) planets (b, c, e) should stay in resonance and expand to the current period ratio $P_{\mathrm{c}}/P_{\mathrm{b}}=2.035$ and $P_{\mathrm{e}}/P_{\mathrm{c}}=3.228$; ii) when planets (b, c, e) reach the current period ratio, planet d should also match the observed period ratio with $P_{\mathrm{d}}/P_{\mathrm{c}}=1.765$ and $P_{\mathrm{e}}/P_{\mathrm{d}}=1.829$. A successful expansion matching these two conditions, as in panels d and g in Figure 2, is not guaranteed.
@@
Because the orbital expansion phase is not related to the formation process of (b, c, e) 3BR, we used an optimized model to form the (b, c, e) 3BR to simplify the parameter setup as the initial masses of the planets. We initialize five planets (b, c, $\mathrm{d_{1}}$, $\mathrm{d_{2}}$ and e) in the first-order resonance chain consistent with Phase I and adiabatically decrease the masses of $\mathrm{d_{1}}$ and $\mathrm{d_{2}}$ to zero to ensure that the remaining planet b, c and e are still in 3BR close to integer period ratio. Then we add planet d back. The detailed setup of the optimized model is discussed in Appendix B.
@@
Figure 5 shows the outcome of two unsuccessful simulations with the same masses of planets following M1 in Table 2 and different assumptions on the initial value of $P_{\mathrm{d}}/P_{\mathrm{c}}$. Initially, the planets are close to their corresponding 2:1 and 3:1 period ratios, which is a condition for the trapping in the zeroth-order 3BR (see Section 4.2). The dashed lines in Figure 5 indicate different 3BRs characterized by $p$, $q$ and $r$ in Equation (7). Tidal dissipation moves planets along $(2,3,5)$ zeroth-order 3BR between b, c, and e (Charalambous et al., 2018; Papaloizou et al., 2018), indicated in the left panels of Figure 5 by the blue dashed line.
@@
While the (b, c, e) system starts from near-exact resonance locations, there is some freedom to choose the initial starting point of planet d. In Figure 5, panels a and b present a simulation where planet d is positioned below the $(p,q,r)=(3,5,8)$ 3BR resonance line. In addition, only tidal damping operates on the planets and all planets have their $Q_{\mathrm{phys}}=4.6$, which is consistent with the successful simulation in Figure 2. These low Q-values are solely adopted for computational expediency since they speed up the simulation. However, as is discussed in more detail in Section 6.3 the young age demands a very efficient damping mechanism. The adopted value for $Q_{\mathrm{phys}}$ in the orbital expansion phase simulations does not affect the simulation outcome. A larger $Q_{\mathrm{phys}}$ will not change the simulation outcome, but it will make the duration of the orbital expansion to the observed period ratio longer. Therefore, the age of the system (younger than $650\;\mathrm{Myr}$) might be inconsistent with the simulation, as explained in Section 6.3. During tidal expansion the periods of planets b, c, and e change, following (by definition) the $(2,3,5)$ 3BR line (panel a), while $P_{\mathrm{d}}$ remains largely constant. As a result, the increase of $P_{\mathrm{e}}$ and decrease of $P_{\mathrm{c}}$ lead to the increase of both $P_{\mathrm{e}}/P_{\mathrm{d}}$ and $P_{\mathrm{d}}/P_{\mathrm{c}}$ and the system moves away from the (3,5,8) resonance line (panel b). Therefore, during the orbital expansion planets c, d, and e would not encounter this 3BR, which could dislodge it from the (b, c, e) resonance (see below), and planets b, c, and e successfully expand to the observed period ratio (panel a). Nevertheless, this scenario fails as the present location of planet d with respect to c an e (the black cross in Figure 5b) cannot be reached.
@@
The slope in the $P_{\mathrm{e}}/P_{\mathrm{d}}$ versus $P_{\mathrm{d}}/P_{\mathrm{c}}$ plane (Figure 5b) is the result of the conservation of the 3BR resonance angle and angular momentum. Specifically, without considering the potentially disturbing effects of planet d, the slope in this plane can be analytically solved (see Appendix C):
$a_{\mathrm{cde}}=\frac{\Delta({P_{\mathrm{e}}/P_{\mathrm{d}}})}{\Delta(P_{\mathrm{d}}/P_{\mathrm{c}})}=\frac{4.90m_{\mathrm{b}}+4.94m_{\mathrm{c}}}{7.27m_{\mathrm{e}}-m_{\mathrm{b}}},$ (10)
@@
A possible solution to match the low $a_{\mathrm{cde}}$ is to decrease the initial semi-major axis of planet d, and thereby $P_{\mathrm{d}}/P_{\mathrm{c}}$, in such a way that it could reach the correct observed position. One example is shown in panels c and d of Figure 5. However, this simulation is also unsuccessful, because planets c, d, and e would encounter the $(3,5,8)$ 3BR during the orbital expansion, as shown in panel d. The resonance overlapping experienced at this point would break the b, c, and e 3BR and stop the orbital expansion of these planets (Petit et al., 2020). We find that the (b, c, e) 3BR are dislodged in more than 95% simulations, similar to what happened in panel c.
@@
    A (temporary) slowdown in the expansion of $P_{\mathrm{d}}/P_{\mathrm{c}}$ due to outward migration of planet c, for example through a positive torque acting on planet b. Because planets b, c, and e are linked by 3BR, the outward migration of b would also lead to the increase in the semi-major axis of c and e. An example of a scenario that includes positive torques is shown in panels a and b of Figure 6. Specifically, a positive torque, with $t_{\mathrm{a0,IV}}^{\mathrm{b}}=2.1\times 10^{10}\,\mathrm{yr}$ and $t_{\Gamma}^{\mathrm{b}}=1.4\,\mathrm{Gyr}$, is added to planet b (see below, Section 5.2.1). The torque slows down the expansion of $P_{\mathrm{d}}/P_{\mathrm{c}}$ with respect to $P_{\mathrm{e}}/P_{\mathrm{d}}$. Its decaying nature results in an S-shaped trajectory in the (c, d, e) period ratio plane (Figure 5b). As a result, the planets successfully reach the observed period ratio near the end of the simulation (see Section 5.2.1).
@@
Goldberg & Batygin (2021) expanded b, c, and e 3BR, but planet d was neglected because its current position is far from a 3BR with any other planet. We stress, however, that planet d is crucial in the orbital expansion phase. Although planet d is far from 3BR currently, when planets c, d, and e hit a 3BR during the orbital evolution, it could break the 3BR between b, c, and e and result in a different architecture from the observation. Therefore, it is necessary to include planet d in the orbital expansion simulation. The initial, relative positions of planet d and the (c, d, e) 3BR are crucial for the subsequent orbital expansion of the system.